Calculate the duration of a bond with two years to maturity, 7% coupon, 8.2% yield and pays semiannually.

1.89 years

1.82 years

1.9 years

3.8 years

None of the above
Answer #1
Year Cash Flow PV Factor PV Of Cash Flow Weight of cash flow Weight * Year
a b c=1/1.07^a d=b*c e=d/1021.70
1 $                82.00 0.93458 $                   76.64 0.0750 0.0750
2 $          1,082.00 0.87344 $                 945.06 0.9250 1.8500
Total $             1,021.70 1.000 1.9
Macaulay Bond Duration =1.9 years